# Learning a Unified Degradation-aware Representation Model for Multi-modal Image Fusion

### v2

@@ 3.3.2 Training
$\mathcal{L}_{1}=\mathcal{L}_{\mathrm{task}}+\alpha_{1}^{(1)}\mathcal{L}_{\mathrm{unified}}+\alpha_{2}^{(1)}\mathcal{L}_{\mathrm{recon}}+\alpha_{3}^{(1)}\mathcal{L}_{\mathrm{text}}$ (4)
where,
@@
    $\mathcal{L}_{\mathrm{recon}}=||\hat{y}_{pd}-y||_{1}$ represents the reconstruction loss for the pseudo-degradation task.
@@
    $\mathcal{L}_{\mathrm{task}}=||\hat{y}-y||_{1}+\alpha_{task}|||\nabla\hat{y}|-|\nabla y|||_{1}$ represents the loss for different degradation tasks ($\nabla$: Sobel operator)
@@
    $\alpha_{1}^{(1)},\alpha_{2}^{(1)},\alpha_{3}^{(1)},\alpha_{\mathrm{task}}$ are hyperparameters.
@@
$Z_{\mathrm{pd}}^{\mathrm{ir}}=E_{\mathrm{img}}(x_{ir},c_{\mathrm{pd}},1),\quad Z_{\mathrm{pd}}^{\mathrm{vi}}=E_{\mathrm{img}}(x_{vi},c_{\mathrm{pd}},0)$ (5)
@@
$O_{\mathrm{fused}}=D\{zero[F(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{pd}}^{\mathrm{ir}})]+rule(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{pd}}^{\mathrm{ir}})\}$ (6)
Inspired by ControlNet [59], we first implement a prior fusion rule, $rule(\cdot,\cdot)$, for initial feature fusion. A learnable module, $F(\cdot,\cdot)$, then refines rule-based fusion results. A zero-initialized $1\times 1$ convolution mitigates random initialization noise and accelerates Second Stage convergence.
@@
$\mathcal{L}_{2}=\mathcal{L}_{\mathrm{color}}+\alpha_{1}^{(2)}\mathcal{L}_{\mathrm{grad}}+\alpha_{2}^{(2)}\mathcal{L}_{\mathrm{per}}$ (7)
where $\alpha_{1}^{(2)}$ and $\alpha_{2}^{(2)}$ are hyperparameters balancing loss contributions. Since the proposed unified latent space can extract powerful degradation irrelevant features, theoretically, the fusion loss function not need to be specially designed. Thus, the color consistency item ($\mathcal{L}_{\mathrm{color}}$) and gradient item ($\mathcal{L}_{\mathrm{grad}}$) are chosen from Text-IF. The perceptual item ($\mathcal{L}_{\mathrm{per}}$) is chosen from U2Fusion.
@@ 3.3.3 Inference
$Z_{\mathrm{pd}}^{\mathrm{vi}}=E_{\mathrm{img}}(x_{vi},c_{pd},0),\quad Z_{\mathrm{t}}^{\mathrm{ir}}=E_{\mathrm{img}}(x_{ir},c_{t},1)$ (8)
$O_{\mathrm{fused}}=D\{zero[F(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{t}}^{\mathrm{ir}})]+rule(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{t}}^{\mathrm{ir}})\}$ (9)
Our End-to-end inference uses a single model. In degradation description, non-degraded and degraded modalities are $c_{\textrm{pd}}$ and $c_{t}$, respectively. This enables diverse degradation handling via modality-specific conditions, crucially addressing combined degradations through description combination, even without combined degradation training data.
Notably, our model do not complicates fusion process. Affected modalities are readily identified from user input via a simple text classification or regular expressions, allowing straightforward degradation description.
Figure 5: Qualitative comparison of degradation-aware fusion tasks: Low Light (LL), Dehaze (HZ), Overexposure (OE), Low Contrast (LC), Super Resolution (SR4, SR8), Overexposure for visible images/Low Contrast for infrared images (OE+LC), Low Light for visible images/Low Contrast for infrared images (LL+LC), and Low Contrast (LC). “eir” denotes External Image Restoration Methods. Text-IF uses “eir” only for Super Resolution and combined degradations. For more detail about the qualitative comparison, please refer to supplementary materials.
@@ 4.1 Setting
In the proposed model, the encoder and the decoder both consist of 4 layers. The channel numbers for each layer are [48, 96, 192, 384], reduced to [16, 32, 64, 128] for each encoder layer’s output size. $K_{bt}$ and $K_{tb}$ are set to [1,1,2,2] and [2,2,4,8], respectively. Decoder blocks and fusion blocks for each layer is set to [1,1,1,1] and [1, 1, 2, 2]. The number of attention heads for TGA is set to 4, and the prior rule is selected from DenseFuse [22].  [^1] More detailed hyperparameter (e.g. hyperparameter ablation) and training settings are provided in the supplementary materials.
@@ 4.1.1 Training Data
For the first stage of training, we utilize the most common tasks for infrared modality: low contrast (LC) and low resolution (SR) ($\times$4 and $\times$8). We use the EMS Full dataset [56] and construct a high-quality super-resolution dataset on LLVIP [15], following the approach in InstructIR [36].
For visible modality, we perform dehazing (HZ), overexposure correction (OE), and low-light enhancement (LL). The high-quality images are chosen from the RESIDE [18], Exposure-Errors [1], and LOL datasets [51], respectively.
For the second stage, the training data of the MSRS datasets [41] is used. All of the aforementioned datasets are publicly available.
All task texts are generated by LLaMA [44].
@@ 4.1.2 Evaluation Data
For vanilla image fusion, we evaluate the fusion performance on the RoadScene [54], M3FD [41], TNO [43], and MSRS [41] test dataset.
For degradation-aware image fusion, we use EMS-full [56] for low contrast, overexposure correction, and low-light enhancement. We also construct a super-resolution test dataset on M3FD [41] in the same manner. For dehazing, we collect hazy samples from M3FD [41] to create a real-world dehazing test dataset.
@@ 4.1.3 Evaluation Metrics
The evaluation metrics include Correlation Coefficient (CC) [12], Spatial Correlation of Differences (SC) [4], Structural Similarity Index (SS) [49], Peak Signal-to-Noise Ratio (PS), Multiscale Structural Similarity (MS) [50], Average Gradient (AG), Standard Deviation (SD), Spatial Frequency (SF), Entropy (EN) [39], CLIP-IQA (CL) [45], LIQE (LI) [60], MANIQA (MA) [55], NIQE (NI) [35], NUSIQ [17] (NU), and ARNIQA (AR) [2].
These metrics comprehensively evaluate the fused image quality from multiple perspectives, including image noise, texture details, and preservation of source image structure. CLIP-IQA [45], LIQE [45], MANIQA [55], NUSIQ [17], and ARNIQA [2] are state-of-the-art no-reference image quality assessment metrics. Except for NIQE, higher values indicate better fusion quality.
@@ 4.2 SOTA Competitors
We conducted comparisons against seven state-of-the-art (SOTA) methods, including DCINN (DCI) [46], FusionBooster (FuB) [8], CSCFuse (CSC) [65], MMDRFuse (MMD) [11], Text-IF (TIF) [56], DAFusion (DAF) [48], and GTMFuse (GTM) [34]. Among these, Text-IF and DAFusion are categorized as All-in-One degradation-aware image fusion models.
@@ 4.3 Vanilla Image Fusion
Table 1: Quantitative comparison of vanilla image fusion task on MSRS and M3FD. Bold/underlined values: best/second best results.
MSRS [41] M3FD [31]
Methods CC SC PS SS MS CC SC PS SS MS
DCI’23 0.58 1.49 32.54 0.36 0.45 0.59 1.37 31.97 0.49 0.49
CSC’23 0.53 0.92 31.83 0.16 0.28 0.58 1.67 30.47 0.43 0.48
MMD’24 0.61 1.53 32.47 0.48 0.52 0.62 1.40 31.13 0.49 0.50
GTM’24 0.60 1.61 31.64 0.42 0.50 0.59 1.52 30.34 0.24 0.41
FuB’25 0.67 1.53 31.2 0.34 0.50 0.59 1.53 29.61 0.40 0.47
TIF’24 0.60 1.68 32.29 0.48 0.52 0.58 1.41 31.67 0.48 0.48
DAF’25 0.66 1.58 28.80 0.26 0.41 0.63 1.75 30.51 0.42 0.46
LURE 0.61 1.74 32.56 0.49 0.53 0.64 1.68 32.12 0.50 0.53
Table 2: Quantitative comparison of vanilla image fusion task on TNO and RoadScene. Bold/underlined values: best/second best results.
TNO [43] RoadScene [54]
Methods CC SC PS SS MS CC SC PS SS MS
DCI’23 0.42 1.61 31.81 0.29 0.36 0.56 1.16 31.49 0.16 0.36
CSC’23 0.44 1.72 30.49 0.44 0.45 0.63 1.80 30.82 0.46 0.52
MMD’24 0.46 1.59 31.18 0.48 0.43 0.63 1.10 31.27 0.41 0.48
GTM’24 0.45 1.68 30.46 0.37 0.41 0.60 1.55 29.55 0.33 0.45
FuB’25 0.46 1.54 29.07 0.40 0.42 0.59 1.03 29.39 0.35 0.45
TIF’24 0.42 1.67 31.29 0.47 0.44 0.61 1.57 31.21 0.48 0.52
DAF’25 0.49 1.70 30.53 0.44 0.41 0.64 1.65 31.34 0.48 0.53
LURE 0.46 1.84 31.92 0.51 0.47 0.57 1.67 31.52 0.50 0.55
@@ 4.4 Degradation-aware Image Fusion
Table 3: Quantitative comparison of degradation-Aware fusion tasks: Low Light Enhancement (LL), Dehaze (HZ), Overexposure Correction (OE), and Low Contrast Enhancement (LC). Bold/underlined values: best/second best results. $\dagger$ denotes All-in-One type methods.
LL HZ OE LC
Method AG SD CL CL LI AR CL LI MA CL MA NI $\downarrow$
DCI’23 4.74 40.90 0.19 0.21 1.17 0.52 0.19 1.26 0.22 0.14 0.14 4.74
CSC’23 4.27 51.08 0.17 0.25 1.09 0.51 0.19 1.38 0.24 0.12 0.12 6.38
MMD’24 3.95 36.07 0.15 0.23 1.17 0.53 0.19 1.23 0.24 0.14 0.14 4.55
GTM’24 4.15 50.75 0.12 0.18 1.01 0.36 0.14 1.05 0.16 0.14 0.13 6.54
FuB’25 4.10 37.36 0.17 0.24 1.12 0.53 0.13 1.05 0.20 0.15 0.15 4.66
TIF’24 $\dagger$ 4.05 42.10 0.19 0.25 1.17 0.53 0.23 1.49 0.23 0.16 0.13 4.41
DAF’25 $\dagger$ 5.19 47.99 0.19 0.25 1.19 0.51 0.20 1.55 0.24 0.13 0.10 3.17
LURE $\dagger$ 4.77 55.23 0.20 0.30 1.14 0.54 0.21 1.37 0.25 0.17 0.14 4.67
Table 4: Quantitative comparison of degradation-aware fusion tasks: Super Resolution (SR4, SR8), Overexposure for visible images/Low Contrast for infrared images (OE+LC), Low Light for visible images/Low Contrast for infrared images (LL+LC). Bold/underlined values: best/second best results. $\dagger$ denotes All-in-One type methods.
@@
Method SF CL AR SF CL AR SD NU CL EN SD CL
DCI’23 10.09 0.26 0.58 10.15 0.24 0.56 42.46 5.85 0.20 7.10 42.46 0.13
CSC’23 14.23 0.28 0.59 13.16 0.25 0.57 47.61 4.20 0.15 5.88 47.61 0.11
MMD’24 9.22 0.29 0.59 9.36 0.25 0.57 37.44 4.98 0.21 6.93 37.44 0.10
GTM’24 8.76 0.32 0.37 8.65 0.29 0.35 53.00 6.36 0.15 7.23 53.00 0.10
FuB’25 9.13 0.28 0.58 9.16 0.24 0.57 34.39 5.57 0.16 6.88 34.39 0.11
TIF’24 $\dagger$ 13.20 0.26 0.59 12.78 0.23 0.57 49.43 6.86 0.23 7.26 49.43 0.12
DAF’25 $\dagger$ 15.23 0.22 0.57 15.23 0.24 0.58 52.56 6.37 0.20 7.24 50.84 0.11
LURE $\dagger$ 11.20 0.33 0.60 9.77 0.27 0.59 55.40 6.84 0.24 7.42 55.40 0.15
@@
For fair comparison, we use a two-step strategy for degradation-unware methods: SOTA external image restoration followed by image fusion. For Text-IF’s limitations on some tasks, e.g., super-resolution and combined degradation, we also use this two-stage approach on these tasks. In addition, DCINN [46], FusionBooster [46], CSCFuse [65], MMDRFuse [11], and GTMFuse [34] all employ a two-step strategy across tasks. Task-specific SOTA image restoration models: URetinex (low-light) [52], SGID-PFF (dehazing) [5], CoTF (overexposure) [29], AirNet (low contrast) [19], SwinFuSR (super-resolution) [3].  [^3] For more information including prompts used in experiments, please refer to supplementary materials..

### v3

@@ 3.3.2 Training
$\mathcal{L}_{1}=\mathcal{L}_{\mathrm{task}}+\alpha_{1}^{(1)}\mathcal{L}_{\mathrm{unified}}+\alpha_{2}^{(1)}\mathcal{L}_{\mathrm{recon}}+\alpha_{3}^{(1)}\mathcal{L}_{\mathrm{text}},$ (4)
where
@@
[…]
@@
[…]
@@
    $\alpha_{1}^{(1)},\alpha_{2}^{(1)},\alpha_{3}^{(1)}$ are hyperparameters.
@@
$Z_{\mathrm{pd}}^{\mathrm{ir}}=E_{\mathrm{img}}(x_{ir},c_{\mathrm{pd}},1),\quad Z_{\mathrm{pd}}^{\mathrm{vi}}=E_{\mathrm{img}}(x_{vi},c_{\mathrm{pd}},0).$ (5)
@@
$O_{\mathrm{fused}}=D\{zero[F(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{pd}}^{\mathrm{ir}})]+rule(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{pd}}^{\mathrm{ir}})\}.$ (6)
Inspired by ControlNet [zhang2023adding], we first implement a prior fusion rule, $rule(\cdot,\cdot)$, for initial feature fusion. A learnable module, $F(\cdot,\cdot)$, then refines rule-based fusion results. A zero-initialized $1\times 1$ convolution mitigates random initialization noise and accelerates Second Stage convergence.
@@
$\mathcal{L}_{2}=\mathcal{L}_{\mathrm{color}}+\alpha_{1}^{(2)}\mathcal{L}_{\mathrm{grad}}+\alpha_{2}^{(2)}\mathcal{L}_{\mathrm{per}},$ (7)
where $\alpha_{1}^{(2)}$ and $\alpha_{2}^{(2)}$ are hyperparameters balancing loss contributions. Since the proposed unified latent space can extract powerful degradation irrelevant features, theoretically, the fusion loss function not need to be specially designed. Thus, the color consistency item ($\mathcal{L}_{\mathrm{color}}$) and gradient item ($\mathcal{L}_{\mathrm{grad}}$) are chosen from Text-IF [yi2024text]. The perceptual item ($\mathcal{L}_{\mathrm{per}}$) is chosen from U2Fusion [xu2020u2fusion].
@@ 3.3.3 Inference
$Z_{\mathrm{pd}}^{\mathrm{vi}}=E_{\mathrm{img}}(x_{vi},c_{pd},0),\quad Z_{\mathrm{t}}^{\mathrm{ir}}=E_{\mathrm{img}}(x_{ir},c_{t},1),$ (8)
$O_{\mathrm{fused}}=D\{zero[F(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{t}}^{\mathrm{ir}})]+rule(Z_{\mathrm{pd}}^{\mathrm{vi}},Z_{\mathrm{t}}^{\mathrm{ir}})\}.$ (9)
End-to-end inference uses one model, where $c_{\textrm{pd}}$ and $c_{t}$ denote non-degraded and degraded modalities. This design enables diverse and combined degradation handling via descriptive conditions, without requiring joint degradation training data.
Figure 5: Qualitative comparison of degradation-aware fusion tasks. “eir” denotes External Image Restoration Methods. Text-IF uses “eir” only for Super Resolution and combined degradations. For more detail about the qualitative comparison, please refer to supplementary materials.
@@ 4.1 Setting
In the proposed model, the encoder and the decoder both consist of 4 layers. The channel numbers for each layer are [48, 96, 192, 384], reduced to [16, 32, 64, 128] for each encoder layer’s output size. $K_{bt}$ and $K_{tb}$ are set to [1,1,2,2] and [2,2,4,8], respectively. Decoder blocks and fusion blocks for each layer is set to [1,1,1,1] and [1, 1, 2, 2]. The number of attention heads for TGA is set to 4, and the prior rule is selected from DenseFuse [li2018densefuse].  [^1] More detailed hyperparameter (e.g. hyperparameter ablation) and training settings are provided in the supplementary materials
@@ 4.1.1 Training Data
For the first stage of training, we utilize the most common tasks for infrared modality: low contrast (LC) and low resolution (SR) ($\times$4 and $\times$8). We use the EMS Full dataset [yi2024text] and construct a high-quality super-resolution dataset on LLVIP [jia2021llvip], following the approach in InstructIR [oh2024instructir]. For visible modality, we perform dehazing (HZ), overexposure correction (OE), and low-light enhancement (LL). The high-quality images are chosen from the RESIDE [li2019benchmarking], Exposure-Errors [afifi2021learning], and LOL datasets [wei2018deep], respectively. Our method does not entirely exclude synthetic datasets but reduces reliance on low-quality ones, using real-world and high-quality synthetic restoration data.
For the second stage, the training data of the MSRS datasets [Tang2022PIAFusion] is used. All of the aforementioned datasets are publicly available.
All task texts are generated by LLaMA [touvron2023llama].
@@ 4.1.2 Evaluation Data
For vanilla image fusion, we evaluate the fusion performance on the RoadScene [xu2020aaai], M3FD [Tang2022PIAFusion], TNO [tno], and MSRS [Tang2022PIAFusion] test dataset.
For degradation-aware image fusion, we use EMS-full [yi2024text] for low contrast, overexposure correction, and low-light enhancement. We also construct a super-resolution test dataset on M3FD [Tang2022PIAFusion] in the same manner. For dehazing, we collect hazy samples from M3FD [Tang2022PIAFusion] to create a real-world dehazing test dataset.
@@ 4.1.3 Evaluation Metrics
The evaluation metrics include Correlation Coefficient (CC) [deshmukh2010image], Spatial Correlation of Differences (SC) [aslantas2015new], Structural Similarity Index (SS) [wang2002universal], Peak Signal-to-Noise Ratio (PS), Multiscale Structural Similarity (MS) [wang2003multiscale], Average Gradient (AG), Standard Deviation (SD), Spatial Frequency (SF), Entropy (EN) [roberts2008assessment], CLIP-IQA (CL) [wang2023exploring], LIQE (LI) [zhang2023liqe], MANIQA (MA) [yang2022maniqa], NIQE (NI) [mittal2012making], NUSIQ [ke2021musiq] (NU), and ARNIQA (AR) [agnolucci2024arniqa].
These metrics comprehensively evaluate the fused image quality from multiple perspectives, including image noise, texture details, and preservation of source image structure. CLIP-IQA [wang2023exploring], LIQE [wang2023exploring], MANIQA [yang2022maniqa], NUSIQ [ke2021musiq], and ARNIQA [agnolucci2024arniqa] are state-of-the-art no-reference image quality assessment metrics. Except for NIQE, higher values indicate better fusion quality. Red and blue values indicate best and second best results.
@@ 4.2 SOTA Competitors
We conducted comparisons against seven state-of-the-art (SOTA) methods, including LRRNet (LRRN) [li2023lrrnet], TITA (TITA) [hu2025balancing], SAGE (SAGE) [wu2025every], MMDRFuse (MMDR) [deng2024mmdrfuse], Text-IF (Text) [yi2024text], DAFusion (DAFu) [wang2025degradation], and CDDFuse (CDDF) [zhao2023cddfuse]. Among these, Text-IF and DAFusion are categorized as All-in-One degradation-aware image fusion models. It should be noted that DAFusion (DAFu) is not included for fairness, since it inherently performs restoration on all inputs.
@@ 4.3 Vanilla Image Fusion
Table 1: Quantitative comparison of vanilla image fusion task on MSRS and M3FD.
RoadScene Dataset MSRS Dataset
Methods EN SC PS $S$ MS EN SC PS $S$ MS
LRRN’23 7.13 1.57 59.94 0.34 0.49 6.19 0.79 64.74 0.21 0.37
CDDF’23 7.45 1.71 62.14 0.48 0.53 6.70 1.62 64.41 0.50 0.51
MMDR’24 6.51 1.10 62.55 0.41 0.48 6.48 1.53 64.95 0.48 0.52
Text’24 6.64 1.57 62.43 0.48 0.52 6.65 1.68 64.59 0.48 0.52
SAGE’25 7.02 1.52 60.85 0.42 0.51 6.00 1.42 65.28 0.45 0.48
TITA’25 7.09 1.30 63.02 0.49 0.51 6.57 1.61 65.01 0.49 0.52
Ours 7.45 1.67 63.04 0.50 0.55 6.67 1.74 65.12 0.49 0.53
Table 2: Quantitative comparison of vanilla image fusion task on TNO and RoadScene.
TNO Dataset M3FD Dataset
Methods EN SC PS $S$ MS EN SC PS $S$ MS
LRRN’23 6.84 1.62 62.66 0.40 0.43 6.33 1.38 64.14 0.39 0.43
CDDF’23 6.94 1.79 62.43 0.50 0.46 6.75 1.59 62.94 0.52 0.52
MMDR’24 6.31 1.59 62.37 0.48 0.43 6.33 1.40 62.27 0.49 0.50
Text’24 7.04 1.67 62.59 0.47 0.44 6.71 1.41 63.35 0.48 0.48
SAGE’25 6.77 1.79 62.00 0.49 0.45 6.74 1.66 62.69 0.50 0.52
TITA’25 6.70 1.62 62.71 0.51 0.43 6.53 1.33 63.75 0.51 0.50
Ours 6.72 1.84 63.84 0.51 0.47 6.67 1.67 64.24 0.50 0.53
@@ 4.4 Degradation-aware Image Fusion
Table 3: Quantitative comparison of degradation-Aware fusion tasks. $\dagger$ denotes All-in-One type methods.
$L$ HZ OE LC
[…]
LRRN’23 3.72 36.41 0.19 0.29 1.07 0.56 0.19 1.24 0.23 0.14 0.17 5.01
CDDF’23 5.49 50.87 0.18 0.28 1.06 0.54 0.20 1.41 0.25 0.14 0.14 4.25
MMDR’24 3.95 36.07 0.15 0.23 1.17 0.54 0.19 1.23 0.24 0.14 0.14 4.55
SAGE’25 4.73 43.59 0.18 0.28 1.21 0.54 0.20 1.17 0.20 0.15 0.14 4.48
TITA’25 4.93 41.68 0.18 0.29 1.17 0.54 0.18 1.45 0.24 0.14 0.14 6.38
Text’24 $\dagger$ 4.05 42.10 0.19 0.26 1.16 0.54 0.23 1.49 0.23 0.16 0.13 4.41
DAFu’25 $\dagger$ 5.19 47.99 0.19 0.25 1.19 0.52 0.20 1.55 0.24 0.13 0.10 3.17
Ours 4.77 55.23 0.20 0.30 1.14 0.55 0.21 1.37 0.26 0.17 0.15 4.67
Table 4: Quantitative comparison of degradation-aware fusion tasks. $\dagger$ denotes All-in-One type methods.
@@
Methods SF CL AR SF CL AR EN AG CL EN SD CL
LRRN’23 10.60 0.30 0.59 10.25 0.26 0.58 7.02 4.68 0.20 6.19 31.73 0.14
CDDF’23 14.00 0.32 0.59 13.54 0.28 0.58 7.47 6.56 0.22 6.70 43.48 0.14
MMDR’24 9.22 0.29 0.59 9.36 0.25 0.57 7.19 4.98 0.21 6.93 37.44 0.10
SAGE’25 11.81 0.26 0.59 11.74 0.23 0.58 7.30 5.74 0.22 5.99 36.26 0.14
TITA’25 12.96 0.30 0.60 12.59 0.26 0.59 7.43 6.13 0.20 7.06 42.66 0.11
Text’24 $\dagger$ 13.20 0.26 0.59 12.78 0.23 0.57 7.51 6.86 0.23 7.26 49.43 0.12
DAFu’25 $\dagger$ 15.23 0.22 0.57 15.23 0.24 0.58 7.47 6.37 0.20 7.24 50.84 0.11
Ours 11.20 0.33 0.60 9.77 0.27 0.59 7.48 6.88 0.24 7.42 55.40 0.15
@@
For fair comparison, we use a two-step strategy for degradation-unware methods: SOTA external image restoration followed by image fusion. For Text-IF’s limitations on some tasks, e.g., super-resolution and combined degradation, we also use this two-stage approach on these tasks. In addition, LRRNet [li2023lrrnet], TITA [hu2025balancing], SAGE [wu2025every], MMDRFuse [deng2024mmdrfuse], and CDDFuse [zhao2023cddfuse] all employ a two-step strategy across tasks. Task-specific SOTA image restoration models: URetinex (LL) [wu2022uretinex], SGID-PFF (HZ) [bai2022self], CoTF (OE) [li2024real], AirNet (LC) [AirNet], SwinFuSR (SR) [Arnold_2024_CVPR].  [^3] For more information including prompts used in experiments, please refer to supplementary materials.